Problem

A chair of mass 10.0 kg is sitting on the horizontal floor; the floor is not frictionless. You push on the chair with a force F = 45.0 N that is directed at an angleof 36.0 degree below the horizontal and the chair slides along the floor.

1-Use Newton's laws to calculate the normal force that the floor exerts on the chair.
Answer #1
normal force=Fsin36+mg
=45sin36+10x9.81=53.46 N
answered by: Arshad
Answer #2

Givens:
mass = m = 10.0 kg
F = 45.0 N
θ = 36.0 below the horizontal

next stepSince our goal is to determine thenormal force, we are only interested in the vertical (y) components of this question. There are three forces that come into play vertically:

N = normal force F(g) = Force of Gravity F(y) = y-component of your push

N = N F(g) = -mg F(y) = -(F)sin(θ)

next step(it isequal to 0 because there is no ay= 0)

N = 124 Newtons

answered by: Kieosha
Answer #3
mg+fsin(theta)=10*9.8+45*sin36=124.45
answered by: Cody Davidson
Answer #4

let normal reaction be N
N+Fsin(36)=mg
N=mg-Fsin(36)
=10*9.8-45*sin(36)
N =71.54N

answered by: wendi
Answer #5
Normal Force = N = 10 x 9.8 + 45sin(45) = 129.82 N
answered by: Math(: